# Kang Chen♣,♡,∗, Zhihao Liu♢,♡,∗, Tonghe Zhang♯,⊳,∗, Zhen Guo, Si Xu, Hao Lin, Hongzhi Zang, & Xiang Li, Quanlu Zhang, Zhaofei Yu, Guoliang Fan, Tiejun Huang, Yu Wang♠,†, Chao Yu♠,♡,†

### v2

@@ 5.2.1 LIBERO
We perform few-shot SFT following the official training configs provided by openpi. For the $\pi_{0}$ model, we utilized a subset of 58 trajectories, sampled from the total of 1,692 trajectories spanning the four task suites in the official LIBERO SFT dataset [^3] https://huggingface.co/datasets/physical-intelligence/libero, to perform SFT, which served as the initial checkpoint [^4] https://huggingface.co/RLinf/RLinf-Pi0-SFT-Spatial-Object-Goal for subsequent RL training on LIBERO-Spatial, LIBERO-Object and LIBERO-Object task suites. Additionally, a larger pool of 208 trajectories was employed for the LIBERO-Long few-shot SFT [^5] https://huggingface.co/RLinf/RLinf-Pi0-SFT-Long due to the long-horizon and more challenging nature of these tasks. For the $\pi_{0.5}$ model, given its better pretrained checkpoint and training config, we only leveraged 40 trajectories for few-shot SFT, providing a unified checkpoint [^6] https://huggingface.co/RLinf/RLinf-Pi05-SFT across task suites.
@@
RL Procedure. In RL, the VLA model receives a multi-modal input state comprising: an agent-view and a wrist-view (both 224 $\times$ 224 RGB images), natural language guidance, the robot end effector pose, and the gripper state. The model outputs an action to interact with the LIBERO environment, which provides a binary reward of 1 for successful task completion and 0 otherwise.
@@
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$, which fine-tunes the few-shot SFT $\pi_{0}$ and $\pi_{0.5}$ models with our proposed Flow-Noise and Flow-SDE, against several state-of-the-art VLAs trained on the entire LIBERO dataset, including Octo, OpenVLA, OpenVLA-OFT, $\pi_{\text{fast}}$ (pertsch2025fast), $\pi_{0}$, and $\pi_{0.5}$. We conduct experiments on four LIBERO task suites and report performance as the success rate across all 500 initial states (10 sub-tasks $\times$ 50 states each).
Analysis. As detailed in Sec.˜5.1, our proposed two solutions, Flow-Noise and Flow-SDE, not only achieve comparable performance but also establish a new state-of-the-art by significantly boosting the performance of the few-shot $\pi_{0}$ and $\pi_{0.5}$ SFT models.
@@
For the few-shot $\pi_{0}$ model, the SFT baseline performs poorly, with an average success rate of only 57.6%, indicating that the model struggles with limited demonstration data. Our proposed $\pi_{\texttt{RL}}$ substantially boosts performance, with Flow-SDE and Flow-Noise reaching 96.1% and 97.6%, respectively, and surpassing the full-dataset $\pi_{0}$ SFT baseline of 94.2%.
@@
While the $\pi_{0.5}$ few-shot SFT baseline achieves a decent average performance of 77.1%, it struggles with the challenging LIBERO-Long task, scoring only 43.9%. Our proposed $\pi_{\texttt{RL}}$ framework rectifies this deficiency, boosting the LIBERO-Long success rate from 43.9% to 94.0%, constituting a 50.1% improvement. Notably, despite using only a single trajectory for SFT, $\pi_{\texttt{RL}}$ reaches 98.3% final performance, surpassing the 96.9% full-dataset SFT model.
Discussion on two methods. Flow-SDE and Flow-Noise differ primarily in their noise injection strategy and MDP formulation, with experiments indicating that Flow-Noise marginally outperforms Flow-SDE, a result we attribute to two factors:
  • Noise Injection: Flow-Noise employs a noise network for exploration, complemented by a relative entropy bonus for noise magnitude adaptation, which affords the model finer control during convergence, thus achieving better performance.
  • MDP Formulation: Flow-Noise adopts a one-layer MDP formulation where the log-probability of the executed action is derived from the joint log-probability of the denoised sequence. This formulation endows Flow-Noise with higher data utilization efficiency, leading to faster convergence, as demonstrated in Fig.˜8.
Despite this, the performance discrepancy is still marginal (e.g., $1.5\%$ in $\pi_{0}$ and $0.4\%$ in $\pi_{0.5}$). Additionally, Flow-Noise requires recomputing the entire denoising trajectory for log-likelihood computation. Consequently, the update time per RL training step scales with the number of denoising steps, whereas it remains constant for Flow-SDE due to its mixed ODE-SDE rollout strategy.
5.2.2 ManiSkill
SFT Procedure. Since the SFT dataset provided by RL4VLA lacks the state information required for $\pi_{0}$, we re-synthesized trajectories following their setting using the MPLib motion planning suite (Guo_MPlib), with the final 15 additional frames appended to reinforce the concept of completing a motion.
RL Procedure. In RL, the VLA model receives an input comprising a single 480 x 640 RGB third-person view, a short language instruction, and the current joint pose. The model also receives a structured reward signal from the environment: 1.0 for correct object placement and 0.1 for successful attachment of the gripper to the object, mitigating unwanted throwing behaviors.
Experiments. Based on the models $\pi_{0}$ and $\pi_{0.5}$ models, we empirically validate the performance of Flow-SDE and Flow-Noise against SFT baselines in SIMPLER and a pick-and-place generalization benchmark. Following the default settings of $\pi_{0}$ and $\pi_{0.5}$’s official code base, we include proprioception information in the input to the action expert of $\pi_{0}$ and omit the state input to the VLM of $\pi_{0.5}$.
  • In SIMPLER, the experimental setup comprises an 8-DoF WidowX-250S arm evaluated on four standard tasks: (1) Spoon: placing a spoon on a cloth, (2) Carrot: placing a carrot on a plate, (3) Eggplant: placing an eggplant in a basket, and (4) Cube: stacking a cube. For the SFT stage, we employ a curated dataset in which each task is trained with 144 demonstration episodes.
  • In the generalization test, the policy is prompted to pick from 16 different object types and place them onto 17 different receptacles, distributed across 16 unique table scenes, yielding a total of 4,352 unique task combinations. Given the high complexity of this setting, the SFT data set was prepared with 16,384 episodes, a scale substantially larger than that for SIMPLER tasks.
Analysis. As detailed in Tab.˜2 and Tab.˜3, $\pi_{\texttt{RL}}$ achieves substantial performance improvements in both the SIMPLER and generalization environments. In the SIMPLER environment, $\pi_{\texttt{RL}}$ increases the average success rate of the $\pi_{0}$ model from 67.2% to 86.7%, with three tasks (carrot, eggplant, and spoon) exceeding 90% success. In the training environment of the generalization test, which comprises 4352 task compositions, the performance of $\pi_{0}$ increases from 41.6% to 85.7%, while the $\pi_{0.5}$ model improves from 40.1% to 84.8%. These results demonstrate the effectiveness of $\pi_{\texttt{RL}}$ in a photorealistic environment.
Table 2: Evaluation results on the WidowX SIMPLER benchmark for $\pi_{0}$ and $\pi_{0.5}$.
Model SIMPLER
Carrot Eggplant Spoon Cube Avg.
$\pi_{0}$ SFT 82.7 87.5 61.7 37.1 67.2
Flow-Noise 95.7 96.7 91.6 63.0 86.7
$\Delta$ +13.0 +9.2 +29.9 +25.9 +19.5
$\pi_{0.5}$ SFT 70.6 91.9 43.5 31.0 59.2
Flow-Noise 82.0 98.2 82.8 53.3 79.1
$\Delta$ +11.4 +6.3 +39.3 +22.3 +19.9
Table 3: Evaluation results on the Generalization Test of ManiSkill.
Model IND OOD
Vision Semantic Execution Avg.
$\pi_{0}$ SFT 38.4 32.6 8.4 13.2 18.1
Flow-SDE 78.8 61.1 25.4 31.5 39.3
Flow-Noise 77.8 63.4 23.1 24.2 36.9
$\Delta$ +40.4 +30.8 +16.8 +18.3 +21.3
$\pi_{0.5}$ SFT 40.1 40.2 16.6 22.4 26.4
Flow-SDE 90.9 68.0 34.5 45.4 49.3
Flow-Noise 89.7 69.9 35.5 54.9 53.4
$\Delta$ +50.8 +29.7 +18.9 +32.5 +27.1
Generalization Tests. Following RL4VLA, we further evaluate the model’s generalization across three challenging out-of-distribution (OOD) scenarios: (1) Vision, challenging the model with novel backgrounds and textures; (2) Semantics, probing comprehension with unseen objects, varied instructions, and confounding elements like extra objects or receptacles; (3) Execution, assessing robustness against varied initial states, unseen robot poses, and dynamic disturbances, such as moving target object during execution. In the OOD scenarios detailed in Tab.˜3, we observe that the $\pi_{0}$-SFT model demonstrates strong generalization for visual information. This can be attributed to the robust foundation of its VLM, which allows it to better handle visual disturbances.
Furthermore, while RL yields significant improvements on in-distribution tasks, we observe its gains are limited in OOD scenarios. We attribute this discrepancy to two factors we aim to address in future work. First, the SFT baseline model itself exhibits substantial performance degradation in OOD settings, which inherently caps the generalization potential achievable by the subsequent RL finetuning. Second, freezing the VLM during the RL stage for training efficiency prevents the model from adapting its visual features to the environment, consequently hindering its visual generalization capabilities.
5.2.3 MetaWorld
RL Procedure. During the RL procedure, the VLA model processes a multi-modal input comprising a 480 $\times$ 480 RGB agent-view image, language guidance, the robot’s end-effector position, and its gripper state. Based on this input, the model outputs an action to interact with the environment, which in turn provides a sparse reward: 1 for successful task completion and 0 otherwise.
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$ against Diffusion Policy (diffusion_policy), TinyVLA (wen2025tinyvla), and SmolVLA (shukor2025smolvla). For the performance evaluation, we follow the setup from SmolVLA, i.e., classifying 50 tasks into easy, medium, hard, and very hard four categories according to their difficulties.
Analysis. As detailed in Tab.˜4, RL fine-tuning substantially boosts performance. The $\pi_{0}$ and $\pi_{0.5}$ models achieve average success rates of 85.8% and 70.7%, respectively. This marks a significant improvement over their SFT-only counterparts and surpasses the best-performing baseline SmolVLA of 68.2%, confirming that RL can effectively enhance model capabilities across a diverse range of manipulation task types.
@@ 5.2.3 MetaWorld
(c) Spatial
(d) Object
(e) Goal
(f) Long
Figure 5: Visual comparison of PPO and GRPO with Flow-SDE $\pi_{0}$ on the LIBERO, demonstrating that PPO outperforms GRPO in terms of convergence performance and training speed.
5.3 Ablation Study
Given that Flow-SDE achieves performance comparable to Flow-Noise while offering higher computational efficiency, we conduct ablation studies with the Flow-SDE method on the LIBERO benchmark to investigate the impact of the RL algorithm, critic design, stochasticity injection strategy, MDP formulation, and various hyperparameters.
5.3.1 RL algorithms
Given the significant performance gains from PPO on the LIBERO benchmark, we also investigated the effectiveness of GRPO (shao2024grpo) (see Appendix A for a detailed description), another widely used policy gradient method applied in VLA+RL training. We compare the performance of PPO and GRPO on both the $\pi_{0}$ and $\pi_{0.5}$ models, with results summarized in Tab.˜5.
We further visualize training curves of PPO and GRPO in Fig.˜5, demonstrating that PPO outperforms GRPO in both final convergence performance and training stability across all four LIBERO task suites.

### v3

@@ C.1 LIBERO Benchmark
We perform few-shot SFT following the official training configurations. For the $\pi_{0}$ model, we fine-tune on a subset of 58 trajectories sampled from the 1,692 total demonstrations [^3] https://huggingface.co/datasets/physical-intelligence/libero, which serves as the initial checkpoint for subsequent RL training on the Spatial, Object, and Goal task suites. For the Long task suite, a larger pool of 208 trajectories is employed to address its more challenging, long-horizon nature. In contrast, for the $\pi_{0.5}$ model, benefiting from a superior pre-trained checkpoint and training configurations, we leverage only 40 trajectories to provide a unified few-shot SFT checkpoint across all task suites.
@@
RL Procedure. In RL, the VLA model receives a multi-modal input state comprising: an agent-view and a wrist-view, natural language guidance, the robot end effector pose, and the gripper state. The model outputs an action to interact with the LIBERO environment, which provides a binary reward of 1 for successful task completion and 0 otherwise.
@@
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$, which fine-tunes the few-shot SFT $\pi_{0}$ and $\pi_{0.5}$ models with Flow-Noise and Flow-SDE, against several state-of-the-art VLAs trained on the entire LIBERO dataset, including Octo, OpenVLA, OpenVLA-OFT, $\pi_{\text{fast}}$ (Pertsch et al., 2025), $\pi_{0}$, and $\pi_{0.5}$. We conduct experiments on four LIBERO task suites and report performance as the success rate across all 500 initial states (10 sub-tasks $\times$ 50 states each).
@@
Analysis. For the few-shot $\pi_{0}$ model, the SFT baseline performs poorly, with an average success rate of only 57.6%, indicating that the model struggles with limited demonstration data. $\pi_{\texttt{RL}}$ substantially boosts performance, with Flow-SDE and Flow-Noise reaching 96.1% and 97.6%, and surpassing the full-dataset $\pi_{0}$ SFT baseline of 94.2%.
@@
While the $\pi_{0.5}$ few-shot SFT baseline achieves a decent average performance of 77.1%, it struggles with the challenging Long task suite, scoring only 43.9%. Our proposed $\pi_{\texttt{RL}}$ rectifies this deficiency, boosting the Long task success rate from 43.9% to 94.0%, constituting a 50.1% improvement. Notably, despite using only a single trajectory for SFT, $\pi_{\texttt{RL}}$ reaches 98.3% final performance, surpassing the 96.9% full-dataset SFT model.
Table 5: Evaluation results on the CALVIN benchmark (Scene D), reporting the average completed subtasks and success rates for task sequences of length 1 to 5.
Methods CALVIN-D
Len-1 Len-2 Len-3 Len-4 Len-5 Avg. $\Delta$ Avg.
$\pi_{0}$ SFT 94.7 84.9 74.3 65.2 57.5 3.766
Flow-SDE 96.4 88.0 77.5 70.8 61.7 3.944 +0.178
Flow-Noise 96.9 88.8 78.0 68.3 59.9 3.919 +0.153
$\pi_{0.5}$ SFT 92.7 84.3 76.7 68.8 61.3 3.838
Flow-SDE 99.7 98.2 95.8 91.0 87.0 4.717 +0.879
Flow-Noise 99.6 97.6 93.9 89.6 84.5 4.652 +0.814
@@
C.2 ManiSkill Benchmark
SFT Procedure. In the ManiSkill benchmark, the policy is required to pick from 16 object types and place them onto 17 receptacles across 16 unique table scenes, yielding 4,352 unique task combinations. Given the high complexity of this setting, the SFT dataset consists of 16,384 episodes synthesized using the MPLib motion planning suite (Guo et al., 2025). To reinforce the concept of motion completion, 15 additional frames are appended to the end of each trajectory.
RL Procedure. In RL, the VLA model receives a third-person RGB image, a concise language instruction, and the current joint proprioception. The environment provides a structured reward signal: 1.0 for correct object placement and an auxiliary 0.1 reward for successful gripper-object attachment, intended to encourage stable manipulation and mitigate undesired behaviors such as impulsive throwing.
Experiments. Following the RL4VLA experimental protocol, we conduct RL training on a comprehensive set of 4,352 task combinations and record the performance as the aggregate success rate across these tasks.
Analysis. As detailed in Tab.˜4, $\pi_{\texttt{RL}}$ significantly boosts performance in the training environment. Specifically, the success rate of $\pi_{0}$ increases from 38.4% to 77.8%, while $\pi_{0.5}$ improves from 40.1% to 90.9%. These gains underscore the efficacy of RL in complex settings.
C.3 CALVIN Benchmark
SFT Procedure. We conduct SFT on the CALVIN ABC dataset [^4] https://huggingface.co/datasets/InternRobotics/InternData-Calvin_ABC, which comprises approximately 24 hours of unstructured "play data" across three distinct environments (A, B, and C). This dataset includes over 20,000 language-labeled trajectories covering 34 unique manipulation tasks.
RL Procedure. Each episode consists of a sequence of five randomly sampled subtasks to be completed in succession without environment resets. The reward signal is defined at the subtask level, where the model receives a sparse binary reward of 1.0 for each successfully executed subtask and 0.0 otherwise.
Table 7: Specific generalization evaluation results in the ManiSkill OOD setting.
Flow-SDE Flow-Noise Flow-SDE Flow-Noise
In distribution Main-v3-train 38.4 78.8 77.8 40.1 90.9 89.7
Visual-Language Variations Instruct-v1-test 30.1 64.6 66.5 46.6 77.0 85.7
VisionImage-v1-test 38.3 68.8 71.7 46.2 78.8 83.1
VisionTexture03-v1-test 35.1 66.0 66.8 36.7 69.6 75.0
VisionTexture05-v1-test 31.0 55.8 60.5 32.7 58.0 62.2
VisionWhole03-v1-test 35.4 62.4 69.0 40.1 69.6 71.6
VisionWhole05-v1-test 28.5 49.0 53.9 30.7 55.0 57.0
Semantic Reasoning (object/receptacle confounders) MultiCarrot-v1-test 7.8 28.2 23.0 16.7 36.8 38.2
MultiCarrot-v1-train 12.5 36.5 31.8 28.2 49.5 50.1
MultiPlate-v1-test 5.0 16.4 18.3 11.8 29.4 28.3
MultiPlate-v1-train 7.3 20.5 19.6 9.7 22.3 25.4
Action Execution PositionChangeTo-v1-test 9.6 17.4 10.9 13.5 36.2 54.7
Position-v1-test 16.9 45.6 37.5 31.2 54.5 55.0
Experiments. We evaluate the performance of $\pi_{\texttt{RL}}$ in Scene D over 1,000 episodes. Following the standard CALVIN evaluation protocol, we report two key metrics: (1) the success rate for task sequences of increasing lengths, namely Len-1 to Len-5. (2) the average number of completed sub-tasks, denoted as Avg. per episode.
Analysis. As detailed in Tab.˜5, $\pi_{\texttt{RL}}$ yields substantial performance gains in long-horizon sequential execution, particularly with the $\pi_{0.5}$ variant. The SFT models inherently struggle with compounding errors across sequential tasks, with $\pi_{0.5}$ only achieving a 61.3% success rate on Len-5 sequences. RL effectively mitigates this issue with the average completed sub-tasks of Flow-SDE increasing from 3.838 to 4.717, and its Len-5 success rate surges to 87.0%.
Notably, the performance gap between SFT and RL widens significantly as the sequence length increases. For $\pi_{0.5}$, while Flow-SDE shows a modest 7.0% improvement over SFT in Len-1 tasks, the gap expands to an impressive 25.7% in the most challenging Len-5 sequences.
C.4 MetaWorld Benchmark
RL Procedure. During the RL procedure, the VLA model processes a multi-modal input comprising a RGB agent-view image, language guidance, the robot’s end-effector position, and its gripper state. Based on this input, the model outputs an action to interact with the environment, which in turn provides a sparse reward: 1 for successful task completion and 0 otherwise.
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$ against Diffusion Policy (Chi et al., 2025), TinyVLA (Wen et al., 2025), and SmolVLA (Shukor et al., 2025). For the performance evaluation, we follow the setup from SmolVLA, i.e., classifying 50 tasks into easy, medium, hard, and very hard four categories according to their difficulties.
Analysis. As detailed in Tab.˜6, RL fine-tuning substantially boosts performance. The $\pi_{0}$ and $\pi_{0.5}$ models achieve average success rates of 85.8% and 70.7%, respectively. This marks a significant improvement over their SFT-only counterparts and surpasses the SmolVLA baseline of 68.2%, confirming that RL can effectively enhance model capabilities across a diverse range of manipulation task types.
Appendix D Out-of-Distribution RL Evaluation
While previous experiments demonstrate significant RL-driven improvements in the ID domain, this section evaluates OOD generalization. As LIBERO lacks a dedicated interface for OOD testing, we utilize the ManiSkill, CALVIN, and MetaWorld benchmarks to investigate whether the RL-driven improvements represent genuine skill acquisition that scales to novel settings, or merely reflect the exploitation of environment-specific biases.
Table 8: Evaluation results on the SIMPLER benchmark for $\pi_{0}$ and $\pi_{0.5}$ with Flow-Noise method.
Model SIMPLER
Carrot Eggplant Spoon Cube Avg.
$\pi_{0}$ SFT 82.7 87.5 61.7 37.1 67.2
+RL 95.7 96.7 91.6 63.0 86.7
$\Delta$ +13.0 +9.2 +29.9 +25.9 +19.5
$\pi_{0.5}$ SFT 70.6 91.9 43.5 31.0 59.2
+RL 82.0 98.2 82.8 53.3 79.1
$\Delta$ +11.4 +6.3 +39.3 +22.3 +19.9
Figure 12: Real-world deployment of an RL refined policy performing a pick and place task.
D.1 ManiSkill
Setup. Following RL4VLA, we evaluate the model’s generalization across three challenging OOD scenarios: (1) Vision, challenging the model with novel backgrounds and textures; (2) Semantics, probing comprehension with unseen objects, varied instructions, and confounding elements like extra objects or receptacles; (3) Execution, assessing robustness against varied initial states, unseen robot poses, and dynamic disturbances.
Results. In the OOD scenarios detailed in Tabs.˜4 and 7, we observe that the $\pi_{0}$-SFT model demonstrates strong generalization for visual information. This can be attributed to the robust foundation of its VLM, which allows it to handle visual disturbances better.
As for the RL training, although the performance improvements in OOD scenarios are lower than those in IND settings, the proportional improvements achieved are notably comparable. As indicated in Tab.˜4, for the $\pi_{0.5}$ model, Flow-SDE enhances the IND success rate by 126.7%, while the OOD similarly increases by 102.3%. This consistency in relative gains indicates that RL-driven optimization promotes the acquisition of generalized action representations rather than merely overfitting the training environment, thus preserving efficacy under distribution shifts.
Nevertheless, we observe a performance gap in relative improvement between the Vision OOD tasks and the IND domain. Specifically, for $\pi_{0.5}$, the 73.9% gain in visual generalization trails the 126.7% increase observed in the training environment. This discrepancy likely stems from freezing the VLM backbone during the RL stage for computational efficiency, which restricts the model’s ability to adapt its visual grounding features to novel textures and backgrounds.
D.2 CALVIN
Setup. We evaluate environmental and visual OOD robustness based on the ABC $\to$ D protocol in CALVIN. Under this setting, the model is trained on ABC environments and evaluated in a zero-shot manner on Scene D. Scene D introduces significant distribution shifts in terms of visual textures, lighting conditions, and spatial layouts, effectively assessing the agent’s ability to transfer skills to an unfamiliar physical environment.
Results. Under identical $\to D$ training settings, the RL finetuned policy in the ABC environment reaches a 79.1% success rate in the OOD scene D, improving over the 61.3% SFT baseline as shown in Fig.˜5. Aligned with the findings in ManiSkill, this suggests that ID gains can be transferred to OOD settings characterized by visual variations.
D.3 MetaWorld
Setup. We utilize the ML45 benchmark from MetaWorld to evaluate the task-level generalization. This setup consists of 50 distinct robotic manipulation tasks: the agent is trained on 45 base tasks and subsequently evaluated on 5 held-out, unseen tasks, which require the model to generalize its learned manipulation primitives to entirely novel task objectives and workspace configurations.
Results. As evidenced in the Fig.˜5, while success rates show consistent gains within the ID domain, OOD performance is characterized by persistent oscillation throughout the training process. This instability indicates that RL, in its current form, struggles to foster stable cross-category generalization.
Nevertheless, the model retains the OOD skills learned during the SFT phase throughout the RL training process. This highlights a significant advantage over standard SFT, which often causes the model to overfit on expert demonstrations and lose its broader capabilities (Li et al., 2025a). Unlike standard SFT, RL enables ID performance gains while preserving the general knowledge, which indicates that RL provides a more balanced optimization that maintains OOD robustness without catastrophic forgetting.
@@ D.3 MetaWorld
D.4 Summary
In conclusion, our OOD evaluation demonstrates that RL enhances performance for similar tasks but fails to generalize effectively to novel task objectives.
Specifically, RL training effectively enhances robustness against low-level variations such as the visual and execution shifts observed in ManiSkill and CALVIN. This indicates that the model acquires generalized action representations rather than merely overfitting to the training environment.
Regarding high-level generalization on MetaWorld, the model successfully retains the OOD skills inherited from the SFT phase, demonstrating that RL avoids the catastrophic forgetting and overfitting typical of standard imitation learning. However, transferring its performance gains to entirely novel task objectives remains a significant challenge.
Appendix E Case Studies: Single-Task RL Training
While the preceding experiments focused on performance across multi-task benchmarks, this section investigates single-task scenarios where the VLA is trained to master a specific task within a relatively static environment. Specifically, we evaluate our approach on the SIMPLER (Li et al., 2024) and a Real2Sim2Real environment.
E.1 SIMPLER
Setup. In SIMPLER, the experimental setup comprises an 8-DoF WidowX-250S arm evaluated on four standard tasks: (1) Spoon: placing a spoon on a cloth. (2) Carrot: placing a carrot on a plate. (3) Eggplant: placing an eggplant in a basket. (4) Cube: stacking a cube. For the SFT stage, we employ a curated dataset in which each task is trained with 144 demonstration episodes.
Analysis. As detailed in Tab.˜8, $\pi_{\texttt{RL}}$ increases the average success rate of the $\pi_{0}$ model from 67.2% to 86.7%, with three tasks (carrot, eggplant, and spoon) exceeding 90% success.